There are 300 purely competitive farms in the local dairy market. Of the 300 dairy farms, 298 have a cost structure that generates profits of $36 for every $600 invested.



Instructions: Enter your answers as whole numbers.



a. What is the percentage rate of return for these 298 dairies?
percent.



b. The other two dairies have a cost structure that generates profits of $52 for every $400 invested. What is their percentage rate of return?
percent.



c. Assuming that the normal rate of profit in the economy is 10 percent, and firms cannot copy each other's technology, will there be entry or exit? .



Will the change in the number of firms affect the two that earn $52 for every $400 invested? .



What will be the rate of return earned by most firms in the industry in long-run equilibrium?
percent.



If firms can copy each other’s technology, what will be the rate of return eventually earned by all firms?
percent.
1
Expert's answer
2020-04-26T19:05:10-0400

a. What is the percentage rate of return for these 298 dairies?

percent.



"r = \dfrac{\$48}{\$600}\times 100 = 8\%"

b. The other two dairies have a cost structure that generates profits of $52 for every $400 invested. What is their percentage rate of return?



"r = \dfrac{\$52}{\$400}\times 100 = 13\%"

c. Assuming that the normal rate of profit in the economy is 10 percent, and firms cannot copy each other's technology, will there be entry or exit? .


Since the normal profit is 10%, and 298 firms are making a profit of 8%, there will be exit from the industry.


Will the change in the number of firms affect the two that earn $52 for every $400 invested? .


The exit will not affect the rate of return for the two firms.


What will be the rate of return earned by most firms in the industry in long-run equilibrium?


The rate of return earned by most firms is 10%. In the long-run equilibrium, the rate of return is always equal to the normal rate of return.

If firms can copy each other’s technology, what will be the rate of return eventually earned by all firms?


If the firms can copy each other's technology, each firm will eventually earn the normal rate of return, which is 10%.